OpenStudy (anonymous):

help pls... gamma functions.. how $$\frac{3}{2}$$ obtained in $$\Gamma(n+\frac{1}{2})=[n+\frac{1}{2}-1] [n+\frac{1}{2}-2]...........\frac{3}{2}\frac{1}{2}\Gamma(\frac{1}{2})$$ Formula used : $$\Gamma(n+a)=[n+a-1] [n+a-2]......a\Gamma(a)$$ And,how to eliminate $$\frac{3}{2}$$ to get the Ans : $$\frac{1.3.5......(2n-1)\sqrt\pi}{2^n}$$

OpenStudy (rational):

Is \(n\) a positive integer ?

OpenStudy (anonymous):

yes, n is a positive integer

OpenStudy (rational):

then, as you keep subtracting 1, last few terms look like below \[\ldots (2+\frac{1}{2}) (1+\frac{1}{2})(0+\frac{1}{2})\Gamma(\frac{1}{2})\] yes ?

OpenStudy (rational):

basically we're using below recurrence relation \[\Gamma(t+1)=t\Gamma(t)\]

OpenStudy (rational):

\[\Gamma(n+\frac{1}{2}) = (n+\frac{1}{2}-1)\Gamma(n+\frac{1}{2}-1)\]

OpenStudy (anonymous):

n + 1/2 -1 = (2n-1) / 2

OpenStudy (anonymous):

yes , i got it . n= 0, 1,2,3 from last term ..... thank you!!

OpenStudy (anonymous):

Gamma(1/2) = sqrt(pi)

OpenStudy (anonymous):

yes , gamma functions is not defined for n=0

OpenStudy (rational):

okay looks neat :)

OpenStudy (anonymous):

actually you can keep going, for instance Gamma( n + 1/2) = (n + 1/2 -1)...1/2 * -1/2 *-3/2 Gamma(-3/2) But you might as well stop at 1/2, since that is well known value.